# Self-supervised Pre-training for Mapping of Archaeological Stone Wall in Historic Landscapes Using High-Resolution DEM Derivatives

### v1

@@ 3.3 Pseudo Siamese Network
The fused representation $z(x)$ is then passed to a decoder head $Dec(z(x))$ for pixel-wise semantic segmentation of dry-stone walls. Depending on the encoder backbone, we use: 1. U-Net decoder (Ronneberger et al., 2015) for ConNet backbones (e.g., ResNet-50 and Wide ResNet-50), and 2. DPT decoder (Ranftl et al., 2021) for ViT backbones (e.g., ViT-Small/16). This architectural design also benefits from the view-invariant representations learned by the student network during cross-view pre-training. Since the parameters of teacher networks are derived from the student networks trained to align outputs across randomly sampled data views, the fused representations $z(x)$ encode robust, view-invariant structural features transferable across different data views.
@@ 3.4 Implementation
We pre-train three widely used vision backbones: ResNet-50 (He et al., 2016), Wide-ResNet-50-2 (Zagoruyko and Komodakis, 2016), and ViT-S/16 (Dosovitskiy et al., 2021), within the DINO-CV framework. The choice of ResNet-50 and ViT-S/16 is motivated by their comparable model sizes (21M vs. 22M parameters) and contrasting architectural paradigms (convolutional vs. attention-based), enabling a fair comparison of ConvNets and Transformers for the task of dry-stone wall mapping. Wide-ResNet-50-2 is chosen for its stronger learning capacity through a wider architecture, and has demonstrated improved performance across various vision tasks.
Pre-training follows the implementation details of DINO (Caron et al., 2021). ViT-S/16 is optimized using AdamW (Loshchilov and Hutter, 2019), while ResNet-50 and Wide-ResNet-50-2 are trained using SGD with momentum. All models are trained with a batch size of 256 for 100 epochs, distributed over two NVIDIA A100 GPUs. The learning rate is linearly warmed up during the first 10 epochs to an initial value of 0.0005 for ViT-S/16 and 0.001 for the ResNet variants, followed by a cosine decay schedule. The base momentum parameter $m$ for the EMA update (Eq. 2) of the teacher network starts at 0.996 and is gradually increased to 1.0 following a cosine schedule.
We apply data augmentations consistent with those used in DINO (Caron et al., 2021) and BYOL (Grill et al., 2020), including the local and global multi-crop strategy. For ViT-S/16, global crops are randomly sampled from scale ranges $(0.33,1.0)$ and local crops from $(0.1,0.33)$. For ResNet-based models, global crops are randomly sampled from $(0.15,1.0)$ and local crops from $(0.05,0.15)$.
@@
We fine-tune the pre-trained vision backbones for 100 epochs on a labeled dataset (Sec 4.1) for stone wall mapping. ResNet-based models are optimized using SGD with a batch size of 64, an initial learning rate of 0.01, and a weight decay of 0.0001. For the ViT-S/16 backbone, we employ the AdamW optimizer with a batch size of 64, a learning rate of 0.00005, and a weight decay of 0.05. For data augmentation in training, input data is resized into $256^{2}$ using bicubic interpolation, followed by random rotations up to 180 degrees. Image normalization is performed according to ImageNet statistics (Deng et al., 2009). For mask inputs, resizing and rotations are performed using nearest neighbor interpolation to preserve label integrity.
@@ 4.1 Dataset
Table LABEL:tab:datasets provides a comprehensive summary of the key specifics of the datasets used in this study.
@@
For self-supervised pre-training, we apply a dataset called BudjBimLandscape, covering 301 km2 of terrain in the Budj Bim cultural landscape, located in southeastern Australia. The dataset comprises unlabeled aerial imagery and two DEM-derived views: multi-directional hillshading (MHS) and visually anisotropic terrain (VAT), both generated from a high-resolution DEM at 50 cm spatial resolution (see Sec. 3.1). Due to the visual occlusion of low-lying stone walls by dense vegetation in aerial imagery, only the MHS and VAT data views are used for DINO-CV pre-training. The dataset contains a total of 188,006 image tiles of size $400^{2}$ pixels, which are resized to $256^{2}$ for training.
@@
For supervised fine-tuning and model evaluation, we use an annotated dataset called BudjBimArea, covering a 25 km2 region in the northern part of the Budj Bim landscape. This area features one of the highest densities of European historic dry-stone walls in Australia. The region is spatially divided into six equal rectangular zones, as illustrated in Fig. 3. Wall annotations were manually created using the Semi-Automatic Classification Plugin within QGIS (Congedo, 2021). Annotations are provided as binary raster masks in the $xy$-plane, where a pixel value of 1 denotes the presence of a wall, and 0 denotes background (non-wall) regions. The dataset is publicly released to support reproducibility and benchmarking.
The dataset comprises a total of 7,102 image tiles, each with a resolution of $400^{2}$ pixels, which are resized to $256^{2}$ for training and evaluation. We adopt a six-fold leave-one-area-out cross-validation strategy, where in each fold, five of the six areas are used for training and validation, and the remaining area is held out for testing. Within the training subset, data is further split into 80% for training and 20% for validation. This splitting strategy ensures evaluation across diverse spatial regions of dataset and provides a robust assessment of generalization to unseen regions.
@@ 4.2 Results
DINO (Caron et al., 2020) ResNet-50-Siam. 46M ImageNet-1K 75.7 74.3 75.0 60.0
@@
DINO (Caron et al., 2020) ViT-S/16-Siam. 42M ImageNet-1K 76.8 81.1 78.9 65.2
@@
Table 2 reports the supervised stone-wall mapping performance of three backbone architectures: ResNet-50, Wide-ResNet-50-2, and ViT-Small. The models are evaluated under different pre-training strategies: random initialization, supervised pre-training on ImageNet-1K, self-supervised pre-training on ImageNet-1K via DINO, SeCo-100K pre-training using DINO-MC, and our proposed DINO-CV on the BudjBimLandscape dataset. The results demonstrate that DINO-CV uniformly outperforms all other pre-training baselines across all model architectures. Specifically, models pre-trained with DINO-CV achieve the highest F1 scores (ResNet50: 81.2%, Wide ResNet50: 81.4% and ViT-small: 80.1%) and mIoU (ResNet50: 68.3%, Wide ResNet50: 68.6% and ViT-small: 66.8%), indicating superior ability to detect and localize dry-stone walls in complex terrain. These performance gains are evident over both supervised and self-supervised baselines, highlighting the benefit of domain-specific, cross-view self-supervised pre-training enabled by DINO-CV. Importantly, the improvements hold across both ConvNet and ViT backbones, suggesting the generalizability and robustness of the proposed pre-training strategy.
((a))
((b))
((c))
@@
Figure 4 provides a detailed comparison of mIoU scores across six areas of the BudjBimArea dataset for each model backbone and pre-training strategy. Overall, DINO-CV achieves the highest or comparable performance relative to all baseline methods. For both ResNet-50 and Wide ResNet-50-2, DINO-CV shows notable improvements in regions with challenging visual characteristics, such as dense vegetation or ambiguous terrain boundaries (e.g., Area 3 and Area 6). ViT-S/16 also benefits significantly from DINO-CV, particularly in areas where other pre-training methods struggle. For example, in Area 2 and Area 4, DINO-CV with ViT-small backbone improves mIoU by 0.07–0.1 compared to random initialization. These results further support the robustness and generalizability of DINO-CV across diverse spatial regions of dataset and architectural designs.
@@
The results demonstrate that models pre-trained with DINO-CV produce more accurate and spatially coherent segmentation masks compared to those initialized randomly or pre-trained via supervised learning. In Area 3 (Fig. 5–ResNet50), the ResNet-50 model with DINO-CV accurately captures intricate wall structures and subtle boundaries, whereas the randomly initialized baseline fails to detect substantial portions of the wall segments. Similarly, in Area 4 (Fig. 5–WideResNet50), the Wide ResNet-50-2 model with DINO-CV more effectively delineates complex wall boundaries in vegetated terrain, outperforming both the randomly initialized and supervised baselines, which misclassify several stone wall fragments. ViT-S/16 also shows notable qualitative improvements in Area 1 (Fig. 5–ViT-S/16), where DINO-CV enhances the integrity of parallel wall segmentation, compared to the fragmented predictions of the other baselines. These qualitative results align with the quantitative performance trends and highlight the effectiveness of DINO-CV in improving the mapping capacity of both ConvNet and ViT backbones across diverse and challenging landscape settings.
Generalization to Unseen Region
Figure 6: Generalization performance of the DINO-CV model with a ResNet-50-Siamese backbone on an unseen region within the Budj Bim Cultural Landscape. The upper panel shows the aerial imagery, with the blue rectangle marking the training area and the green rectangle denoting the unseen test area. The lower panel zooms into the unseen region, where the model accurately segments stone wall structures (annonated in white lines) beneath dense vegetation. These results highlight DINO-CV’s ability to learn transferable geomorphic representations from multi-view DEM derivatives (MHS and VAT), enabling large-scale mapping beyond the spatial extent of the training data.
To evaluate the spatial generalization of DINO-CV, the fine-tuned ResNet-50-Siamese model was tested on a region (green rectangle) located 10 km from the training area (blue rectangle), as shown in Figs. 6. Despite the absence of labeled samples, the model successfully delineates (in white) linear stone wall structures that are largely obscured by dense vegetation. This demonstrates robust transferability of learned representations across distinct spatial and ecological conditions. The cross-view self-supervision on DEM derivatives (MHS and VAT as inputs) enables the model to capture geomorphic patterns characteristic of archaeological stone walls. These results indicate that DINO-CV provides a scalable framework for large-scale, automated mapping of stone wall networks in complex and vegetated landscapes, reducing reliance on extensive labeled data, and supporting reliable mapping in areas that are inaccessible or hazardous for field survey.
@@ 5 Discussion
Table 3: Ablation study on the DINO-CV. We analyze the effectiveness of two key design components: (1) the effect of cross-view self-supervised pre-training of DINO-CV, and (2) the use of the Pseudo Siamese Architecture for supervised fine-tuning. Each row reports the F1 score and mean Intersection over Union (mIoU) for a specific configuration, evaluated on the BudjBimArea dataset. The columns $\Delta_{F1}$ and $\Delta_{mIoU}$ denote the performance difference relative to the DINO-CV baseline (row 1, 5, and 9 for each backbone).
@@
In Table 3, we conduct an ablation study to evaluate the effectiveness of two key design choices of our mapping approach: (1) the use of DINO-CV for pre-training, and (2) the Pseudo Siamese architecture for fusing representations from multiple DEM derivatives (MHS and VAT) during supervised fine-tuning.
Across all backbones (i.e., ResNet-50, WRN-50-2, and ViT-S/16), removing the cross-view component (rows 2, 6, 10) leads to consistent performance drops in both F1 and mIoU, confirming the effectiveness of cross-view training in learning view-invariant representations. This supports our core hypothesis that encouraging alignment across different DEM modalities enhances generalization and robustness. Furthermore, removing the Pseudo Siamese structure and instead using a single view (either MHS or VAT; rows 3–4, 7–8, 11–12) also results in lower performance compared to the dual-view configuration. The improvements from data view fusion demonstrate the complementary nature of MHS and VAT features, validating our architectural design for dual-stream representation learning. Together, these results show that both the cross-view pre-training and the dual-stream Pseudo Siamese architecture contribute positively and independently to model performance, with the full DINO-CV configuration achieving the best results across all backbones.
@@
To gain insight into the spatial attention patterns of models trained with DINO-CV, we visualize the self-attention maps produced by the ViT-S/16 backbone in Fig. 7, without any supervised fine-tuning. These maps highlight the regions of the input DEM derivatives (i.e., MHS and VAT) that the model attends to during inference. The visualizations reveal that DINO-CV enables the model to consistently focus on linear features resembling dry-stone walls. In particular, attention heads in intermediate ViT layers (i.e., attention layers 4 and 5) emphasize wall-like structures, while other attention layers (i.e., attention layer 1, 3 and 6) capture broader contextual information, such as the surrounding terrain and slope.
To further evaluate the alignment of feature representations across data views, we analyze the cosine similarity between the latent features extracted from MHS and VAT inputs using the teacher networks of DINO-CV. This analysis is conducted on both ViT and ResNet backbones, without any supervised fine-tuning. As shown in Fig. 8, DINO-CV exhibits substantially higher cosine similarity along the diagonal of the similarity matrix of data views compared to the original DINO (without cross-view learning), indicating improved feature alignment and enhanced view-invariance. These results confirm that DINO-CV effectively reduces inter-view feature discrepancies—particularly in mid-level spatial regions—and support the hypothesis that cross-view pre-training enables robust and consistent representations across diverse DEM derivatives.

### v3

@@ 3.4 Pseudo-Siamese Network
The fused representation $z(x)$ is then passed to a decoder head $Dec(z(x))$ for pixel-wise semantic segmentation of dry-stone walls. Depending on the encoder backbone, we use: 1. U-Net decoder (Ronneberger et al., 2015) for ConvNet backbones (e.g., ResNet-50 and Wide ResNet-50), and 2. DPT decoder (Ranftl et al., 2021) for ViT backbones (e.g., ViT-S/16). This architectural design also benefits from the view-invariant representations learned by the student network during cross-view pre-training. Since the parameters of teacher networks are derived from the student networks trained to align outputs across randomly sampled data views, the fused representations $z(x)$ encode robust, view-invariant structural features transferable across different data views.
@@ 3.5 Implementation
We pre-train three widely used vision backbones: ResNet-50 (He et al., 2016), Wide-ResNet-50-2 (Zagoruyko and Komodakis, 2016), and ViT-S/16 (Dosovitskiy et al., 2021), using the DINO-CV framework. The choice of ResNet-50 and ViT-S/16 is motivated by their comparable model sizes (21M vs. 22M parameters) and contrasting architectural paradigms (convolutional vs. attention-based), enabling a fair comparison of different vision backbones for the dry-stone wall mapping. Wide-ResNet-50-2 is selected for its increased representational capacity through a wider architecture, which has demonstrated strong performance across diverse vision tasks.
Pre-training follows the established implementation protocols of DINO (Caron et al., 2021). ViT-S/16 is optimized using AdamW (Loshchilov and Hutter, 2019), while ResNet-50 and Wide-ResNet-50-2 are trained using SGD with momentum. All models are trained with a batch size of 256 for 100 epochs, distributed over two NVIDIA A100 GPUs. The learning rate is linearly warmed up during the first 10 epochs to an initial value of 0.0005 for ViT-S/16 and 0.001 for the ResNet variants, followed by a cosine decay schedule. The base momentum parameter $m$ for the EMA update (Eq. 2) of the teacher network is initialized at 0.996 and gradually increased to 1.0 following a cosine schedule.
We apply data augmentations consistent with DINO (Caron et al., 2021) and BYOL (Grill et al., 2020), including the multi-crop strategy with both global and local views. For ViT-S/16, global crops are randomly sampled from scale ranges $(0.33,1.0)$ and local crops from $(0.1,0.33)$. For ResNet-based models, global crops are randomly sampled from $(0.15,1.0)$ and local crops from $(0.05,0.15)$.
@@
We fine-tune the pre-trained vision backbones for 100 epochs on the labeled BudjBimArea dataset (Sec. 4.1) for stone wall segmentation. Hyperparameters for fine-tuning are selected based on a combination of prior work on DINO (Caron et al., 2021) and empirical validation for training stability under limited labeled data. To maintain compatibility with the learned representations, we retain the same optimizers used during self-supervised pre-training (AdamW for ViT-S/16 and SGD for ResNet-based backbones).
ResNet-based models are optimized using SGD with a batch size of 64, an initial learning rate of 0.01, and a weight decay of 0.0001. For ViT-S/16, we employ the AdamW optimizer with a batch size of 64, a learning rate of 0.00005, and a weight decay of 0.05. The reduced batch size compared to pre-training reflects the increased memory footprint of the pseudo-Siamese architecture (Fig. 3b), which processes MHS and VAT inputs in parallel. For ViT-S/16, a small learning rate is used to preserve the geometric and semantic representations learned during self-supervised pre-training, as larger learning rates were observed to cause unstable convergence and feature degradation. In contrast, ResNet-based models optimized with SGD require a comparatively larger learning rate to achieve effective fine-tuning.
For data augmentation in fine-tuning, input data are resized to $256\times 256$ using bicubic interpolation, followed by random rotations up to 180 degrees. Image normalization is performed according to ImageNet statistics (Deng et al., 2009). For mask inputs, resizing and rotations are performed using nearest-neighbor interpolation to preserve label integrity.
@@ 4 Experiments
Figure 4: Oblique 3D visualization of the annotated airborne LiDAR point clouds within the BudjBimArea dataset. The dry-stone wall structures are highlighted in light yellow (Classification ID 13), demonstrating their visibility beneath the dense vegetation canopy. Despite significant occlusion from trees and undergrowth, the high density of the LiDAR survey allows for the continuous delineation of these low-relief features. These walls typically measure approximately 0.5-1.0 m in width and 0.5-1.2 m in height, as derived from the ground point measurements.
The experimental setting in this study evaluates intra-site generalization (i.e., spatial generalization within a single landscape) under a self-supervised learning paradigm. Self-supervised pre-training is conducted on the BudjBimLandscape dataset using only unlabeled DEM-derived inputs, while supervised fine-tuning and evaluation are restricted to labeled samples from the spatially distinct BudjBimArea subset. This setting reflects a common and practically relevant scenario in archaeological remote sensing, where extensive LiDAR coverage of a site is available but only a limited fraction can be annotated. The objective is therefore to maximize label efficiency and intra-site generalization within a single landscape.
@@ 4.1 Dataset
Table 1 provides a comprehensive summary of the key characteristics of the datasets used in this study.
@@
For self-supervised pre-training, we use the BudjBimLandscape dataset, which covers approximately 301 km2 of terrain within the Budj Bim Cultural Landscape in southeastern Australia. The dataset comprises unlabeled aerial imagery and two DEM-derived visualization products: MHS and VAT, both generated from a high-resolution LiDAR-derived DEM at 10 cm spatial resolution (see Sec. 3.2). Due to the substantial visual occlusion of low-lying stone walls by dense vegetation in optical imagery, only the MHS and VAT representations are used during DINO-CV self-supervised pre-training. The dataset contains a total of 188,006 image tiles of size $400\times 400$ pixels, which are resized to $256\times 256$ for training.
@@
For supervised fine-tuning and evaluation, we use the annotated BudjBimArea dataset, a geographical subset of the larger BudjBimLandscape, covering a 25 km2 region in the northern part of the Budj Bim landscape. This area contains one of the highest densities of historic European dry-stone walls in Australia. The region is spatially partitioned into six non-overlapping rectangular zones, as illustrated in Fig. 5. The dataset comprises 7,102 image tiles, each with a resolution of $400\times 400$ pixels, resized to $256\times 256$ for fine-tuning and evaluation. We adopt a six-fold leave-one-area-out cross-validation strategy, where in each fold five areas are used for training and the remaining area is held out for testing. Within the training subset, data is further split into 80% for training and 20% for validation. This strategy ensures that every tile is evaluated exactly once and provides a robust assessment of intra-site generalization across spatially distinct regions.
Data Annotations
To ensure the reliability of the ground truth labels, a multi-modal annotation strategy is adopted to integrate raster-based labeling with geometric verification using LiDAR data. Dry-stone wall annotations were manually delineated as binary raster masks in the $xy$-plane using the Semi-Automatic Classification Plugin within QGIS (Congedo, 2021). While individual DEM derivatives (i.e., MHS and VAT) might exhibit visual noise due to vegetation cover and terrain roughness, cross-referencing these data views with high-resolution aerial imagery (Fig. 2) enabled the reliable identification of linear structures of dry-stone walls and their distinction from natural topographic features. To further reduce false positives associated with elevated vegetation that may overlap wall footprints in 2D raster views, we incorporated geometric verification using the underlying 3D LiDAR point clouds (Fig.4). In practice, a height-above-ground filter was applied during label assignment. Height above ground was estimated using a local ground surface model derived from neighboring ground points, and points exceeding 2 m above the estimated ground surface were excluded. This threshold reflects the typical vertical extent of dry-stone walls in the study area (approximately 0.5–1.2 m in height). This annotation workflow substantially reduces high-canopy noise and improves label consistency, ensuring that the resulting binary masks accurately represent the spatial extent of archaeological stone wall features.
@@ 4.2 Results
DINO (Caron et al., 2021) ResNet-50-Siam. 46M ImageNet-1K 75.7 74.3 75.0 60.0
@@
DINO (Caron et al., 2021) ViT-S/16-Siam. 42M ImageNet-1K 76.8 81.1 78.9 65.2
@@
Table 2 reports the supervised stone wall mapping performance on three backbone architectures: ResNet-50, Wide-ResNet-50-2, and ViT-S/16. The models are evaluated under different pre-training strategies: random weight initialization, supervised pre-training on ImageNet-1K, self-supervised pre-training on ImageNet-1K via DINO, SeCo-100K pre-training using DINO-MC, and our proposed DINO-CV on the BudjBimLandscape dataset. The results demonstrate that DINO-CV uniformly outperforms all other pre-training baselines across all model architectures. Explicitly, models pre-trained with DINO-CV achieve the highest F1 scores (ResNet50: 81.2%, Wide ResNet50: 81.4% and ViT-S/16: 80.1%) and mIoU (ResNet50: 68.3%, Wide ResNet50: 68.6% and ViT-S/16: 66.8%), indicating strong capability to detect and localize dry-stone walls in complex terrain. These performance gains are evident over both supervised and self-supervised baselines, highlighting the benefit of domain-specific, cross-view self-supervised pre-training enabled by DINO-CV. Importantly, the improvements hold across both ConvNet and ViT backbones, suggesting the strong model-agnostic generalization and robustness of the proposed pre-training strategy.
@@
Figure 6 provides a detailed comparison of mIoU scores across six areas of the BudjBimArea dataset for each model backbone and pre-training strategy. Overall, DINO-CV achieves the highest or comparable performance relative to all baseline methods. For both ResNet-50 and Wide ResNet-50-2, DINO-CV shows notable improvements in regions with challenging visual characteristics, such as dense vegetation or ambiguous terrain boundaries (e.g., Area 3 and Area 6). ViT-S/16 also benefits significantly from DINO-CV, particularly in areas where other pre-training methods struggle. For example, in Area 2 and Area 4, DINO-CV with ViT-S/16 backbone improves mIoU by 7–10 percentage points compared to random initialization. These results further support the robustness and intra-site generalization of DINO-CV across diverse spatial regions of the dataset and architectural designs.
@@
The results demonstrate that models pre-trained with DINO-CV produce more accurate and spatially coherent segmentation masks compared to those initialized randomly or pre-trained via supervised learning. In Area 3 (Fig. 7, ResNet-50), the ResNet-50 model with DINO-CV accurately captures intricate wall structures and subtle boundaries, whereas the randomly initialized baseline fails to detect substantial portions of the wall segments. Similarly, in Area 4 (Fig. 7, Wide ResNet50), the Wide ResNet-50-2 model with DINO-CV more effectively delineates complex wall boundaries in vegetated terrain, outperforming both the randomly initialized and supervised baselines, which misclassify several stone wall fragments. ViT-S/16 also shows notable qualitative improvements in Area 1 (Fig. 7, ViT-S/16), where DINO-CV enhances the integrity of parallel wall segmentation, compared to the fragmented predictions of the other baselines. These qualitative results align with the quantitative performance trends and highlight the effectiveness of DINO-CV in improving the mapping capacity of both ConvNet and ViT backbones across diverse and challenging landscape settings.
Generalization to Unlabeled Region
Figure 8: Intra-site generalization performance of the DINO-CV model with a ResNet-50-Siamese backbone on an unlabeled region within the Budj Bim Cultural Landscape. The upper panel shows the aerial imagery, with the blue rectangle indicating the annotated training area and the green rectangle denoting the spatially separated unlabeled region. The lower panel provides a zoomed view of the unlabeled region, where the model delineates linear dry-stone wall structures (shown in white) beneath dense vegetation.
To qualitatively assess the spatial generalization of DINO-CV beyond the annotated region, the fine-tuned ResNet-50-Siamese model was applied to an unlabeled area (green rectangle) located approximately 10 km from the training area (blue rectangle), as shown in Fig. 8. Despite the absence of labeled samples, the model successfully delineates linear dry-stone wall structures (shown in white) that are largely obscured by dense vegetation. This qualitative result suggests that cross-view self-supervised pre-training on DEM derivatives (MHS and VAT) enables the model to learn transferable geomorphic and geometric latent features that remain effective across spatially separated regions within the same landscape. While this experiment does not constitute independent quantitative validation, it provides supporting evidence for the robustness and scalability of the proposed approach in unlabeled and logistically inaccessible regions.
@@ 5 Discussion
Table 3: Ablation study on the DINO-CV. We analyze the effectiveness of two key design components: (1) the effect of cross-view self-supervised pre-training of DINO-CV, and (2) the use of the Pseudo-Siamese Architecture for supervised fine-tuning. Each row reports the F1 score and mean Intersection over Union (mIoU) for a specific configuration, evaluated on the BudjBimArea dataset. The columns $\Delta_{F1}$ and $\Delta_{mIoU}$ denote the performance difference relative to the DINO-CV baseline (row 1, 5, and 9 for each backbone).
@@
In Table 3, we conduct an ablation study to evaluate the effectiveness of two key design choices of our mapping approach: (1) the use of DINO-CV for pre-training, and (2) the Pseudo-Siamese architecture for fusing representations from multiple DEM derivatives (MHS and VAT) during supervised fine-tuning.
Across all backbones (i.e., ResNet-50, WRN-50-2, and ViT-S/16), removing the cross-view pre-training component (rows 2, 6, 10) leads to consistent performance drops in both F1 and mIoU, confirming the effectiveness of cross-view pre-training in learning view-invariant representations. This finding supports our core hypothesis that enforcing alignment across complementary DEM modalities enhances the model capability for mapping. Furthermore, replacing the dual-stream Pseudo-Siamese architecture with a single-stream architecture (either MHS or VAT; rows 3–4, 7–8, 11–12) also results in lower performance. The improvements brought by the dual-stream data view fusion demonstrate the complementary nature of MHS and VAT features, validating the Pseudo-Siamese architectural design. Together, these results show that both the cross-view pre-training and the dual-stream Pseudo-Siamese architecture contribute positively and independently to model performance, with the full DINO-CV configuration achieving the best results across all backbones.
@@
To gain insight into the spatial attention patterns of models trained with DINO-CV, we visualize the self-attention maps produced by the ViT-S/16 backbone in Fig. 9, without any supervised fine-tuning. These maps highlight the regions of the input DEM derivatives (i.e., MHS and VAT) that the model attends to during inference. The visualizations reveal that DINO-CV enables the model to consistently focus on linear and associated geomorphic features resembling dry-stone walls. In particular, attention heads in intermediate ViT layers (i.e., attention layers 4 and 5) emphasize wall-like structures, while other attention layers (i.e., attention layer 1, 3 and 6) capture broader contextual information, such as the surrounding terrain and slope.
To further evaluate the alignment of feature representations across data views, we analyze the cosine similarity between the latent features extracted from MHS and VAT inputs using the teacher networks of DINO-CV. This analysis is conducted on both ViT and ResNet backbones, without any supervised fine-tuning. As shown in Fig. 10, DINO-CV exhibits substantially higher cosine similarity along the diagonal of the similarity matrix of data views compared to the original DINO (without cross-view learning), indicating improved feature alignment and enhanced view-invariance. These results confirm that DINO-CV effectively reduces inter-view feature discrepancies, particularly in mid-level spatial regions, and support the hypothesis that cross-view pre-training enables robust and consistent representations across diverse DEM derivatives.